Eileen is 20 years old. Her grandmother is 65 years old. How many years ago was the grandmother six times as old as Eileen?
Answer by josh_jordan(263)

Eileen: 20
Grandmother = 65

Years ago = x

6(20 - x) = 65 - x

120 - 6x = 65 - x

55 = 5x

x = 11

Therefore, 11 years ago, the grandmother was six times as old as Eileen.